OpenStudy (nani200):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = 3x2 - 1

OpenStudy (nani200):

I think its even but im not sure

OpenStudy (3mar):

May I help?

OpenStudy (nani200):

ok thanks

OpenStudy (nani200):

yes please

OpenStudy (3mar):

Thank you. With my pleasure! You are in the right track: If it's even, then f(-x) = f(x) If it's odd then f(-x) = -f(x)

OpenStudy (nani200):

so you do 3(-x)^2 -1 ?

OpenStudy (3mar):

\[\Large f(x) = 3x^2 - 1\] \[\Large f(\color{red}{-x}) = 3(\color{red}{-x})^2 - 1=-3x^2-1\neq f(x)\neq -f(x)\]

OpenStudy (3mar):

Does it make sense for you now?

OpenStudy (3mar):

Thank you for the medal! That is so kind of you! @satellite73

OpenStudy (nani200):

yes so its odd because its not the same as it was before

OpenStudy (3mar):

If it's even, then f(-x) = f(x) If it's odd then f(-x) = -f(x) If f(-x) is neither equal to f(x) nor -f(x), so it is ....?

OpenStudy (nani200):

neither?

OpenStudy (3mar):

Yes .... \[\Huge\color{MediumSpringGreen }\checkmark\] You hit the right target!

OpenStudy (nani200):

thank you!

OpenStudy (3mar):

Wait Wait Wait Wait Wait

OpenStudy (3mar):

@nani200

OpenStudy (nani200):

yes?

OpenStudy (3mar):

Why did you let me do this wrong and you did not even comment? \[\Large f(\color{red}{-x}) = 3(\color{red}{-x})^2 - 1=-3x^2-1\]

OpenStudy (nani200):

i didnt notice

OpenStudy (nani200):

what's wrong about it?

OpenStudy (nani200):

@3mar i dont see anything wrong with it

OpenStudy (3mar):

\[\Large f(\color{red}{-x}) = 3(\color{red}{-x})^2 - 1=3x^2-1=f(x)\]